### v1

@@ 4.3.5 Link Prediction
Table 10: Experimental statistics of UMLS, Wikidata5M, and FB15k-237N datasets under the LP subtask in transductive setting.
Model Year UMLS Wikidata5M FB15k-237N MR Hits@10 MR MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 TransE 92 2013 1.84 98.9 109370 25.3 17 31.1 39.2 25.5 15.2 30.1 45.9 TransH 94 2014 1.8 99.5 - - - - - - - - - DistMult 34 2014 5.52 84.6 - 25.3 20.9 27.8 33.4 20.9 14.3 23.4 33 TransR 96 2015 1.81 99.4 - - - - - - - - - TransD 91 2015 1.71 99.3 - - - - - - - - - ComplEx 25 2016 2.59 96.7 - 30.8 25.5 - 39.8 24.9 18 27.6 38 DKRL 36 2016 - - 31566 16 12 18.1 22.9 - - - - ConvE 29 2018 1.51 99 - - - - - 27.3 19.2 30.5 42.9 RoBERTa 83 2019 - - 1381597 0.1 0 0.1 0.3 - - - - RotatE 84 2019 - - - 29 23.4 32.2 39 27.9 17.7 32 48.1 QuatE 78 2019 - - - 27.6 22.7 30.1 35.9 - - - - CompGCN 24 2020 - - - - - - - 31.6 23.1 34.9 48 NePTuNe 69 2021 - - - - - - - - - - - ComplEx-N3-RP 27 2021 - 99.8 - - - - - - - - - KG-BERT 48 2019 1.47 99 - - - - - 20.3 13.9 20.1 40.3 MTL-KGC 66 2020 - - - - - - - 24.1 16 28.4 43 StAR 88 2021 1.49 99.1 - - - - - - - - - MEM-KGC 63 2021 - - - - - - - - - - - LaSS 57 2021 1.56 98.9 - - - - - - - - - KEPLER 47 2021 - - 200267 21 17.3 22.4 27.7 - - - - SimKGC 86 2021 - - - 35.8 31.3 37.6 44.1 - - - - LP-BERT 62 2022 1.18 100 - - - - - - - - - KGT5 52 2022 - - - 33.6 28.6 36.2 42.6 - - - - GenKGC 40 2022 - - - - - - - - 18.7 27.3 33.7 KG-S2S 51 2022 - - - - - - - 35.4 28.5 38.8 49.3 CSPromp-KG 32 2023 - - - 38 34.3 39.9 44.6 36 28.1 39.5 51.1 ReSKGC 81 2023 - - - 39.6 37.3 41.3 43.7 - - - - CD 20 2024 - - - - - - - 37.2 28.8 41 53
@@
Table 11: Experimental statistics under the LP subtask in inductive setting.
Model Year NELL-One Wikidata5M N-Shot MRR Hits@1 Hits@5 Hits@10 N-Shot MRR MR Hits@1 Hits@3 Hits@10 DKRL 36 2016 - - - - - Zero-Shot 23.1 78 5.9 32 54.6 GMatching 41 2018 Five-Shot 20 14 26 31 - - - - - - MetaR 64 2019 Five-Shot 26 17 35 44 - - - - - - GMatching 41 2018 One-Shot 19 12 26 31 - - - - - - RoBERTa 83 2019 - - - - - Zero-Shot 7.4 723 0.7 1 19.6 MetaR 64 2019 One-Shot 25 17 34 40 - - - - - - StAR 88 2021 Zero-Shot 26 17 35 45 - - - - - - MEM-KGC 63 2021 - - - - - - - - - - - KEPLER 47 2021 - - - - - Zero-Shot 40.2 28 22.2 51.4 73 SimKGC 86 2021 - - - - - Zero-Shot 71.4 - 50.9 78.5 91.7 KG-S2S 51 2022 Zero-Shot 31 22 41 48 - - - - - -
@@ 5 Future Directions
The integration of large language models (LLMs) into knowledge representation learning (KRL) has demonstrated significant potential in enhancing the capabilities of artificial intelligence systems. However, there are numerous avenues for further research and development. This section outlines six promising directions to advance the field.
5.1 Advanced Integration of LLMs and KRL for Enhanced Contextual Understanding
One major area for future exploration is the advanced integration of LLMs with KRL to achieve deeper contextual understanding. Current models have shown that LLMs can significantly enhance the representation and reasoning capabilities of knowledge graphs (KGs). However, most existing approaches treat textual and structural data separately or integrate them at a superficial level. Future research could focus on more sophisticated methods of merging these data types.
For instance, developing models that can dynamically switch between textual and structural contexts based on the task requirements could lead to more robust and contextually aware systems. This might involve hybrid architectures that combine the strengths of LLMs for textual data with graph neural networks (GNNs) for structural data. These models could leverage attention mechanisms to weigh the importance of textual and structural information dynamically.
@@ 5.2 Development of Efficient Training and Inference Techniques
The computational demands of LLMs are a significant challenge, particularly when integrating them with KRL tasks. Future research should focus on developing more efficient training and inference techniques to make these models more accessible and practical for real-world applications.
Techniques such as model pruning, quantization, and knowledge distillation have shown promise in reducing the computational load of LLMs without significantly sacrificing performance. Applying these techniques specifically to the context of KRL could yield models that are both powerful and efficient. For example, distilling the knowledge from a large LLM into a smaller model that is optimized for KRL tasks could provide a balance between performance and efficiency.
Another promising direction is the use of transfer learning and continual learning approaches. Instead of training models from scratch for each new task or domain, leveraging pre-trained models and fine-tuning them on specific KRL tasks can save significant computational resources. Continual learning techniques, where models incrementally learn from new data without forgetting previously acquired knowledge, can also enhance the adaptability and efficiency of LLM-based KRL systems.
@@ 5.3 Improving Generalization and Robustness in KRL Models
While LLM-enhanced KRL models have achieved impressive results on benchmark datasets, their generalization to unseen data and robustness in the face of noisy or incomplete data remain critical challenges. Future research should aim to develop models that can generalize better across diverse domains and handle the inherent uncertainty and noise in real-world data.
@@ 5.4 Exploring Interpretability and Explainability in KRL Models
As LLM-enhanced KRL models become more complex, their interpretability and explainability become increasingly important, especially in critical applications such as healthcare, finance, and legal domains. Future research should focus on developing methods to make these models more transparent and understandable.
One approach is to develop techniques that can provide insights into the decision-making process of LLM-based KRL models. This could involve the use of attention mechanisms that highlight the most relevant parts of the input data, as well as the development of visualization tools that can illustrate how different parts of the model contribute to the final predictions.
@@ 5.5 Enhancing Human-AI Collaboration in KRL
The integration of LLMs with KRL offers significant potential for enhancing human-AI collaboration. Future research should explore how these advanced models can be leveraged to support and augment human decision-making and knowledge discovery processes.
Developing interactive systems that allow humans to query and interact with LLM-enhanced KRL models can lead to more effective knowledge exploration and problem-solving. For example, creating natural language interfaces that enable users to ask questions and receive explanations or suggestions from the model can make complex knowledge graphs more accessible and useful.
@@
5.6 Addressing Ethical and Societal Implications
As LLM-enhanced KRL models become more powerful and pervasive, it is crucial to address the ethical and societal implications of their deployment. Future research should focus on ensuring that these models are developed and used responsibly, with considerations for fairness, accountability, and transparency.
One important aspect is to ensure that these models do not perpetuate or amplify biases present in the training data. Techniques such as bias detection and mitigation, as well as the development of fairness-aware algorithms, can help address these issues. It is also important to consider the broader societal impacts of deploying these models, including potential job displacement and the impact on privacy.
Additionally, research should explore the development of regulatory frameworks and ethical guidelines for the use of LLM-enhanced KRL models. This could involve collaborations between researchers, policymakers, and industry stakeholders to ensure that these models are used in ways that align with societal values and ethical principles.
@@ 6 Conclusion
The integration of LLM with KRL represents a transformative advancement in AI, offering significant improvements in the ability to understand, represent, and utilize complex knowledge structures. This survey has systematically reviewed various approaches, categorizing them into encoder-based, encoder-decoder-based, and decoder-based methods, and evaluating their performance across multiple downstream tasks. Finally, there remain critical areas for further research.

### v2

@@ 4.3.5 Link Prediction
Table 10: Experimental statistics of FB15k-237 dataset under the LP subtask in transductive settings.
Model FB15k-237
MR MRR Hits@1 Hits@3 Hits@10
TransE [52] 2013 223 27.9 19.8 37.6 47.4
TransH [76] 2014 255 - - - 48.6
DistMult [71] 2014 411 28.1 199 30.1 44.6
TransR [65] 2015 237 - - - 51.1
TransD [77] 2015 246 - - - 48.4
ComplEx [72] 2016 508 27.8 19.4 29.7 45
ConvE [54] 2018 245 31.2 22.5 34.1 49.7
ConvKB [78] 2018 257 24.3 - - 51.7
R-GCN [79] 2018 600 16,4 10 18.1 41.7
KBGAN [80] 2018 - 27.7 - - 45.8
RotatE [73] 2019 177 33.8 24.1 37.5 53.3
KBAT [81] 2019 270 15.7 - - 33.1
CapsE [82] 2019 403 15 - - 35.6
QuatE [83] 2019 176 31.1 22.1 34.2 49.5
TuckER [84] 2019 - 35.8 26.6 39.4 54.4
HAKE [85] 2019 - 34.6 25 38.1 54.2
AttH [86] 2020 - 34.8 25.2 38.4 54
REFE [87] 2020 - - - - 54.1
GAATs [88] 2020 187 - - - 65
Complex-DURA [89] 2020 56 - - - -
DensE [90] 2020 169 34.9 25.6 38.4 53.5
LineaRE [91] 2020 155 35.7 26.4 39.1 54.5
RESCAL-DURA [92] 2020 - 36.8 27.6 - 55
CompGCN [93] 2020 - 35.5 26.4 39 53.5
NePTuNe [94] 2021 - - - - 54.7
ComplEx-N3-RP [95] 2021 - - - - 56.8
ConE [96] 2021 - 34.5 24.7 38.1 54
Rot-Pro [97] 2021 - 34.4 24.6 38.3 54
QuatDE [98] 2021 90 36.5 26.8 40 56.3
NBFNet [99] 2021 - 41.5 32.1 - 59.9
KG-BERT [19] 2019 153 23.7 16.9 26 42.7
MTL-KGC [20] 2020 132 26.7 17.2 29.8 45.8
Pretrain-KGE [100] 2020 - 35 25 38.4 55.4
StAR [25] 2021 1117 29.6 20.5 32.2 48.2
MEM-KGC [101] 2021 - 34.9 26 38.2 52.4
LaSS [74] 2021 108 - - - 53.3
SimKGC [26] 2021 - 33.6 24.9 36.2 51.1
LP-BERT [27] 2022 154 31 22.3 33.6 49
KGT5 [34] 2022 - 34.3 25.2 - 37.7
OpenWorld KGC [102] 2022 - 34.6 25.3 38.1 53.1
LMKE [75] 2022 141 30.6 21.8 33.1 48.4
GenKGC [32] 2022 - - 19.2 35.5 43.9
KG-S2S [35] 2022 - 33.6 25.7 37.3 49.8
kNN-KGE [103] 2023 - 28 37.3 - -
CSPromp-KG [24] 2023 - 35.8 26.9 39.3 53.8
GPT-3.5 2023 - - 23.7 - -
CP-KGC [39] 2024 - 33.8 25.1 36.5 51.6
KICGPT [41] 2024 - 41.2 32.7 44.8 55.4
@@
Table 11: Experimental statistics of UMLS and Wikidata5M datasets under the LP subtask in transductive setting.
Model Year UMLS Wikidata5M
MR Hits@10 MR MRR Hits@1 Hits@3 Hits@10
TransE [52] 2013 1.84 98.9 109370 25.3 17 31.1 39.2
TransH [76] 2014 1.8 99.5 - - - - -
DistMult [71] 2014 5.52 84.6 - 25.3 20.9 27.8 33.4
TransR [65] 2015 1.81 99.4 - - - - -
TransD [77] 2015 1.71 99.3 - - - - -
ComplEx [72] 2016 2.59 96.7 - 30.8 25.5 - 39.8
DKRL [66] 2016 - - 31566 16 12 18.1 22.9
ConvE [54] 2018 1.51 99 - - - - -
RoBERTa [6] 2019 - - 1381597 0.1 0 0.1 0.3
RotatE [73] 2019 - - - 29 23.4 32.2 39
QuatE [83] 2019 - - - 27.6 22.7 30.1 35.9
ComplEx-N3-RP [95] 2021 - 99.8 - - - - -
KG-BERT [19] 2019 1.47 99 - - - - -
StAR [25] 2021 1.49 99.1 - - - - -
LaSS [74] 2021 1.56 98.9 - - - - -
KEPLER [28] 2021 - - 200267 21 17.3 22.4 27.7
SimKGC [26] 2021 - - - 35.8 31.3 37.6 44.1
LP-BERT [27] 2022 1.18 100 - - - - -
KGT5 [34] 2022 - - - 33.6 28.6 36.2 42.6
CSPromp-KG [24] 2023 - - - 38 34.3 39.9 44.6
ReSKGC [104] 2023 - - - 39.6 37.3 41.3 43.7
Table 12: Experimental statistics of FB15k-237N dataset under the LP subtask in transductive setting.
Model FB15k-237N
MRR Hits@1 Hits@3 Hits@10
TransE [52] 2013 25.5 15.2 30.1 45.9
DistMult [71] 2014 20.9 14.3 23.4 33
ComplEx [72] 2016 24.9 18 27.6 38
ConvE [54] 2018 27.3 19.2 30.5 42.9
RotatE [73] 2019 27.9 17.7 32 48.1
CompGCN [93] 2020 31.6 23.1 34.9 48
KG-BERT [19] 2019 20.3 13.9 20.1 40.3
MTL-KGC [20] 2020 24.1 16 28.4 43
GenKGC [32] 2022 - 18.7 27.3 33.7
KG-S2S [35] 2022 35.4 28.5 38.8 49.3
CSPromp-KG [24] 2023 36 28.1 39.5 51.1
CD [38] 2024 37.2 28.8 41 53
Table 13: Experimental statistics under the LP subtask in inductive setting on NELL-One dataset.
Model Year NELL-One
N-Shot MRR Hits@1 Hits@5 Hits@10
GMatching [55] 2018 Five-Shot 20 14 26 31
GMatching [55] 2018 One-Shot 19 12 26 31
MetaR [105] 2019 Five-Shot 26 17 35 44
MetaR [105] 2019 One-Shot 25 17 34 40
StAR [25] 2021 Zero-Shot 26 17 35 45
KG-S2S [35] 2022 Zero-Shot 31 22 41 48
Table 14: Experimental statistics under the LP subtask in inductive setting on Wikidata5M dataset.
Model Year Wikidata5M
N-Shot MRR MR Hits@1 Hits@3 Hits@10
DKRL [66] 2016 Zero-Shot 23.1 78 5.9 32 54.6
RoBERTa [6] 2019 Zero-Shot 7.4 723 0.7 1 19.6
KEPLER [28] 2021 Zero-Shot 40.2 28 22.2 51.4 73
SimKGC [26] 2021 Zero-Shot 71.4 - 50.9 78.5 91.7
@@ 5 Future Directions
The integration of LLM into KRL has demonstrated significant potential in enhancing the capabilities of artificial intelligence systems. However, there are numerous avenues for further research and development. This section outlines five promising directions to advance the field.
5.1 Advanced Integration of LLM and KRL for Enhanced Contextual Understanding
One major area for future exploration is the advanced integration of LLM with KRL to achieve deeper contextual understanding. Current models have shown that LLM can significantly enhance the representation and reasoning capabilities of knowledge graphs. However, most existing approaches treat textual and structural data separately or integrate them at a superficial level. Future research could focus on more sophisticated methods of merging these data types.
For instance, developing models that can dynamically switch between textual and structural contexts based on the task requirements could lead to more robust and contextually aware systems. This might involve hybrid architectures that combine the strengths of LLM for textual data with graph neural networks for structural data. These models could leverage attention mechanisms to weigh the importance of textual and structural information dynamically.
@@ 5.2 Development of Efficient Training and Inference Techniques
The computational demands of LLM are a significant challenge, particularly when integrating them with KRL tasks. Future research should focus on developing more efficient training and inference techniques to make these models more accessible and practical for real-world applications.
Techniques such as model pruning, quantization, and knowledge distillation have shown promise in reducing the computational load of LLM without significantly sacrificing performance. Applying these techniques specifically to the context of KRL could yield models that are both powerful and efficient. For example, distilling the knowledge from a large LLM into a smaller model that is optimized for KRL tasks could provide a balance between performance and efficiency.
Another promising direction is the use of transfer learning and continual learning approaches. Instead of training models from scratch for each new task or domain, leveraging pre-trained models and fine-tuning them on specific KRL tasks can save significant computational resources. Continual learning techniques, where models incrementally learn from new data without forgetting previously acquired knowledge, can also enhance the adaptability and efficiency of LLM enhanced KRL systems.
@@ 5.3 Improving Generalization and Robustness in KRL Models
While LLM enhanced KRL models have achieved impressive results on benchmark datasets, their generalization to unseen data and robustness in the face of noisy or incomplete data remain critical challenges. Future research should aim to develop models that can generalize better across diverse domains and handle the inherent uncertainty and noise in real-world data.
@@ 5.4 Exploring Interpretability and Explainability in KRL Models
As LLM enhanced KRL models become more complex, their interpretability and explainability become increasingly important, especially in critical applications such as healthcare, finance, and legal domains. Future research should focus on developing methods to make these models more transparent and understandable.
One approach is to develop techniques that can provide insights into the decision-making process of LLM enhanced KRL models. This could involve the use of attention mechanisms that highlight the most relevant parts of the input data, as well as the development of visualization tools that can illustrate how different parts of the model contribute to the final predictions.
@@ 5.5 Enhancing Human-AI Collaboration in KRL
The integration of LLM with KRL offers significant potential for enhancing human-AI collaboration. Future research should explore how these advanced models can be leveraged to support and augment human decision-making and knowledge discovery processes.
Developing interactive systems that allow humans to query and interact with LLM enhanced KRL models can lead to more effective knowledge exploration and problem-solving. For example, creating natural language interfaces that enable users to ask questions and receive explanations or suggestions from the model can make complex knowledge graphs more accessible and useful.
@@ 6 Conclusion
In this paper, we present a comprehensive survey of large language models (LLMs) that enhance Knowledge Representation Learning (KRL). We begin by delineating the utilization of LLMs in Knowledge Graphs (KGs) and summarize various Transformer architecture-based methodologies, including encoder-only, encoder-decoder, and decoder-only frameworks. These classifications facilitate an understanding of the evolution from purely structured representations to hybrid models that integrate structured and textual data. Additionally, we discuss the current research methodologies and their performance benefits in downstream tasks. We also outline available datasets, open-source codebases, and diverse applications. Conclusively, we propose several future research directions that are poised to drive advancements not only in KRL but also across broader fields within Natural Language Processing (NLP), leveraging the robust capabilities of LLMs.